You and a friend play a game with a fair, 6-sided die. If you roll a 1 or a 2, you win $15. If you roll a 3, you lose $20. If you roll a 4 or a 5, you lose $10. If you roll a 6, you win $30. What is your expected gain or loss each time you play? Show your work.
Answer by Boreal(15235)

it is (1/3)(+15)-(1/6)(20)-(1/3)(10)+(1/6)(30)=5-10/3-10/3+5
=10-20/3
=+10/3 or +$3.33